Calculate the pressure, in atmospheres, of 2.40 moles helium gas in a 8.00 L container at 18 °C.
1
Expert's answer
2020-10-19T14:20:48-0400

According to the ideal gas law, "pV=nRT," where p is the pressure, V is the volume, n is amount in moles, T is the temperature in Kelvin, and R is a gas constant that equals "0.082057 L\cdot{atm\cdot{mol^{-1}\cdot{K^{-1}}}}".

"T = 18+273.15=291.15K"

Therefore,


"p=\frac{nRT}{V}=\frac{2.40mol\times0.082057 L\cdot{atm\cdot{mol^{-1}\cdot{K^{-1}}}}\times{291.15K}}{8.00L}=7.17atm"


Answer: 7.17 atm